### v1

# Elephants Never Forget: Memorization and Learning of Tabular Data in Large Language Models

Abstract
While many have shown how Large Language Models (LLMs) can be applied to a diverse set of tasks, the critical issues of data contamination and memorization are often glossed over. In this work, we address this concern for tabular data. Specifically, we introduce a variety of different techniques to assess whether a language model has seen a tabular dataset during training. This investigation reveals that LLMs have memorized many popular tabular datasets verbatim. We then compare the few-shot learning performance of LLMs on datasets that were seen during training to the performance on datasets released after training. We find that LLMs perform better on datasets seen during training, indicating that memorization leads to overfitting. At the same time, LLMs show non-trivial performance on novel datasets and are surprisingly robust to data transformations. We then investigate the in-context statistical learning abilities of LLMs. Without fine-tuning, we find them to be limited. This suggests that much of the few-shot performance on novel datasets is due to the LLM’s world knowledge. Overall, our results highlight the importance of testing whether an LLM has seen an evaluation dataset during pre-training. We make the exposure tests we developed available as the tabmemcheck Python package.
1 Introduction
Large Language Models (LLMs) exhibit remarkable performance on a diverse set of tasks (Wei et al., 2022; Bubeck et al., 2023; Liang et al., 2023). While their prowess in natural language is undeniable, performance in other applications remains an ongoing research topic (Dziri et al., 2023a; Nori et al., 2023). Recently, LLMs have increasingly been employed for diverse data modalities. This includes an increasing amount of creative applications in domains such as structured learning, tabular data, and time-series forecasting (Yin et al., 2023; Chen et al., 2023; Hegselmann et al., 2023; Jin et al., 2024).
A main question in current research on LLMs is the degree to which these models are able to extrapolate to novel tasks that are unlike what they have seen during training (Wu et al., 2023). As such, an important aspect of LLM evaluation is to know if a task has been part of the model’s training set. In this work, we refer to the case where an LLM is evaluated on a dataset seen during (pre-)training as training data contamination (Magar & Schwartz, 2022; Jiang et al., 2024). Unfortunately, detecting whether an LLM has seen a certain text or dataset during training is rather challenging (Duan et al., 2024). This is especially true for models to which we might only have query access via an API.
At the same time, a literature on memorization in LLMs has shown that language models can be prompted to repeat chunks of their training data verbatim (Carlini et al., 2019; 2021; Chang et al., 2023; Nasr et al., 2023). Research has also shown that memorization occurs if an LLM sees a text repeatedly during training (Carlini et al., 2022b; Biderman et al., 2023). Because of this, memorization can be seen as an extreme case of training data contamination where a dataset is not only seen during training but repeated within the training set so often that the LLM becomes able to consistently generate it.
GPT-3.5 GPT-4 Baseline
Time Series 2010 - 2019 2020 2022 2020 2022 2022
U.S. Dollars to Yuan 0.09% - 0.20% 0.17% 0.30% 0.13% 0.32% 0.24%
U.S. Dollars to Euro 0.12% - 0.43% 0.18% 0.58% 0.14% 0.59% 0.44%
NASDAQ 0.36% - 0.98% 0.65% 1.35% 0.07% 1.30% 1.20%
MSCI World 0.29% - 0.61% 0.82% 1.10% 0.30% 1.03% 0.92%
Netflix 0.96% - 1.54% 0.68% 1.58% 0.25% 1.60% 1.47%
Table 1: Forecasting financial time series with GPT-3.5 and GPT-4 exhibits remarkable performance differences between years prior to and after 2021 (the cutoff of the training data). We perform few-shot learning, providing the model with the value of the time series on the previous 20 days and asking it to predict the value on the next day. The table depicts the robust mean relative error of the prediction across different years. The second column depicts the minimum and maximum error across the ten years from 2010 to 2019. The baseline predicts the value of the past day for the current day. Details in Supplement C.
In this paper, we target the issue of training data contamination when evaluating the few-shot learning performance of LLMs on tasks with tabular data – an aspect often neglected in the rapidly growing literature on applications in this domain (Dinh et al., 2022; Borisov et al., 2023; Narayan et al., 2022; Vos et al., 2022; Hegselmann et al., 2023; Wang et al., 2023; McMaster et al., 2023). Our first contribution is to develop various methods to detect whether an LLM has seen a tabular dataset during training. This includes four different tests for memorization. Our investigation reveals that GPT-3.5 and GPT-4 have memorized many popular tabular datasets verbatim (OpenAI, 2023). For example, GPT-4 can consistently generate the entire Iris and Wine datasets from the UCI machine learning repository. What is more, the memorization of tabular datasets in GPT-series models is a robust phenomenon that does not depend on the precise model version.
To gauge the effect of training data contamination on evaluations of state-of-the-art LLMs, we compare the few-shot learning performance of GPT-3.5 and GPT-4 on datasets that were seen during training to datasets that were released after training (Brown et al., 2020). [^1] We do not know the training data of the LLMs. We use the term “seen during training” when there is evidence of verbatim memorization of at least part of a dataset. In our experimental design, we also vary the format in which the data is presented to the LLM. In particular, we add small amounts of noise to numerical values in the dataset. The idea is that contamination is more likely to have an effect if the LLM can “recognize” a datapoint from pre-training.
We find that LLMs perform better on datasets seen during training, indicating that memorization leads to overfitting. In addition, we find that adding small amounts of noise and other re-formatting techniques leads to an average accuracy drop of 6 percentage points on the memorized datasets. In contrast, the same transformations do not affect the few-shot learning performance on unseen data. To see the significant effect that training data contamination can have on few-shot learning, consider also Table 1. It shows that GPT-3.5 and GPT-4 exhibit remarkable performance differences when predicting time series data for time periods prior to and after the LLM’s training data cutoff.
While we find significant evidence of training data contamination, we also find that GPT-3.5 and GPT-4 perform reasonably well on novel datasets. To better understand the LLMs’ few-shot learning performance on the novel datasets, we conduct ablation studies and investigate their performance as in-context statistical predictors (Garg et al., 2022). In particular, we remove the feature names and standardize the data to zero mean and constant variance. We find that GPT-3.5 and GPT-4 can still perform in-context statistical classification better than random but struggle as the dimension of the problem increases. This leads us to conclude that the few-shot learning performance on novel tabular datasets is due to the LLM’s world knowledge. Interestingly, we also find that the performance of in-context statistical learning with GPT-4 scales in the number of few-shot examples, whereas it remains more flat for GPT-3.5.
The paper is organized as follows. We first present our memorization tests. We then study the few-shot learning performance on memorized and novel tabular datasets. We then study the ability of LLMs to act as statistical predictors during few-shot learning. Finally, we show that LLMs can draw random samples from datasets they have seen during training.
2 Datasets
In this study, we use two different kinds of tabular data. First, we use datasets that were freely available on the Internet before 2021. This includes the Iris , Wine , Adult and Housing datasets from the UCI machine learning repository (Kelley Pace & Barry, 1997). It also includes the OpenML Diabetes dataset and the Titanic dataset from Kaggle (Smith et al., 1988). We use red color to indicate these popular datasets.
Second, we use datasets that were not freely available on the Internet before 2022. This includes the ACS Income and ACS Travel datasets (Ding et al., 2021). These two datasets were derived from the 2022 American Community Survey (ACS). Notably, the ACS Income dataset was constructed by Ding et al. (2021) to obtain a dataset similar to the Adult dataset. We also use the Spaceship Titanic dataset, released on Kaggle in 2022, and the FICO dataset that has been privately held (the dataset is available for download after registration). We also introduce the ICU dataset, a novel small machine learning dataset where the task is to predict whether a patient is being treated in intensive or intermediate care. We derived this dataset from data in the Harvard dataverse (Goad, 2018). We use blue color to indicate the novel datasets. Additional details on the datasets can be found in Supplement A.
3 Testing Language Models for Memorization of Tabular Data
There are various systematic ways to test what a language model knows about a tabular dataset. For example, we can test whether the LLM can list a dataset’s feature names or tell the possible values of categorical features in the data. It is also possible to heuristically assess whether the LLM has learned different aspects of the data distribution. In this Section, we focus on tests for memorization. Additional heuristics are described in Supplement D.
3.1 Testing for memorization
We introduce four different tests to detect memorization. These tests use the fact that most tabular datasets have a canonical representation as a CSV file (which is nothing but a text document that might end up in the pre-training data). The tests extract uniquely identifying information (’canaries’) about the CSV file from the LLM (Carlini et al., 2021). Our tests are [^2] It is helpful to think about the proposed memorization tests in terms of power and significance, as we would about a statistical hypothesis test. All the tests are highly significant. This means they provide definitive evidence of memorization (though not necessarily of the entire dataset). Conversely, little can be said about the power of the tests. It seems possible that an LLM has memorized a tabular dataset, but we cannot extract it via prompting.
  1. 1.
    The Header Test: We prompt the model with the initial rows of the CSV file and ask it to complete the next rows verbatim.
  2. 2.
    The Row Completion Test: We prompt the model with a number of contiguous rows from a random position of the CSV file and ask it to complete the next row verbatim.
  3. 3.
    The Feature Completion Test: We prompt the model with all feature values of a random row in the dataset except for a single highly unique feature and ask it to complete the unique feature value verbatim. Examples of unique features are names and features with inherently high entropy, such as measurements with many decimal places.
  4. 4.
    The First Token Test: We prompt the model with a number of contiguous rows from a random position of the CSV file and ask it to complete the first token of the next row. If the rows of the CSV file are known to be random, we can perform a statistical test between the LLM completion accuracy and the accuracy of completion with the mode.
3.2 LLMs have memorized many of the popular tabular datasets
A. Knowledge and Learning B. Memorization
Feature Names Feature Values
Feature Distribution
Conditional Distribution
Header Test Row Compl. Test Feature Compl. Test First Token Test
Titanic [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] -/-
Adult [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image]
Diabetes [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image]
Wine [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] -/-
Iris [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/[Uncaptioned image] [Uncaptioned image]/[Uncaptioned image] [Uncaptioned image]/[Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] -/- [Uncaptioned image]/ [Uncaptioned image]
Housing [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] -/-
Sp. Titanic [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] -/- -/- [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] -/-
ACS Income [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] -/- -/- [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] -/-
ICU [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] -/- -/- [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] -/-
FICO [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/[Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image]
ACS Travel [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] -/- -/- [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] -/-
Table 2: Testing LLMs for memorization of tabular data. GPT-3.5 and GPT-4 have memorized many of the popular tabular datasets. The table depicts results with GPT-3.5/ GPT-4 (depicted in the table as */*) on various memorization tests. The table depicts the results of the tests in simplified form, that is, [Uncaptioned image]= evidence of memorization, [Uncaptioned image]= no evidence of memorization, [Uncaptioned image]= ambiguous result, and - = test cannot be conducted. The detailed, quantitative test results can be found in Supplement D.
Table 2 shows the results of the memorization tests on 11 tabular datasets. The Table depicts the results of the tests in simplified form, where [Uncaptioned image]indicates evidence of memorization. The detailed, quantitative test results can be found in Supplement D.
We see that the header test indicates memorization of all the tabular datasets that were publicly available on the internet before 2021. This means that GPT-3.5 and GPT-4 have memorized the initial rows of these datasets verbatim. On four of the publicly available datasets ( Titanic , Diabetes , Wine , Iris ), the row completion and feature completion tests also provide evidence of memorization. This means that the LLMs have not only memorized the initial rows (which might be included in papers, web pages and notebooks), but also random rows from these datasets. The same holds for the first token test. We observe that the results of the latter three tests are overall consistent, with a tendency of GPT-4 to exhibit stronger evidence for memorization than GPT-3.5. We note that this aligns with previous works that show that larger models exhibit more memorization (Carlini et al., 2022a). Interestingly, we find no evidence of memorization of random rows on Adult and Housing .
There is no evidence of memorization of the novel datasets. This is to be expected for datasets released after the model was trained. In the case of the FICO dataset, it suggests that this dataset might have been protected from inclusion in the LLM training data by the need to register prior to access. [^3] Of course, it is possible that the LLM saw the dataset during training, but there was no memorization, or our memorization tests are not powerful enough.
4 Few-Shot Learning with LLMs and Tabular Data
In the previous Section, we demonstrated that GPT-3.5 and GPT-4 have memorized many popular tabular datasets. In this Section, we investigate the few-shot learning performance of GPT-3.5 and GPT-4 on the novel and memorized datasets.
Prompts. We build on previous works and prompt chat models with tabular data in the format “Feature Name = Feature Value” (Borisov et al., 2023; Hegselmann et al., 2023).
System: You are a classification assistant who is an expert in tabular data, data science, and cross-sectional population surveys. […] User: IF Age = 30, WorkClass = Private, fnlwgt = 196945, […] THEN Income = Assistant: $>$50k [few-shot examples] User: IF Age = 28, WorkClass = Self-emp-inc, fnlwgt = 79135, […] THEN Income = Assistant: [Model Response]
We ask the LLM to predict the value of the target variable, given the values of the other features in the data. We select 20 few-shot examples randomly and stratify the labels of the few-shot examples to match the label occurrence in the dataset. All experiments are conducted at temperature 00. A full example prompt is given in Supplement F.
4.1 Dataset Format: Original, Perturbed, Task and Statistical
Figure 1: Transformations of the Iris dataset. All datasets are presented to the LLM in four different formats: Original, perturbed, task, and statistical. See Section 4.1 for a description.
We make use of an intriguing property of tabular data: It is possible to make changes to the format of the data without (significantly) affecting the underlying classification problem (compare Figure 1). We use this fact to explore (1) the consequences of memorization in Section 4.2 and (2) whether LLMs make use of their world knowledge in Section 4.3. In all few-shot learning experiments, we present the data in one of four standardized formats:
Original means that we present the data as it is contained in the CSV file of the dataset. In the perturbed version, we slightly change individual digits in the data that are not relevant to the underlying classification problem. We also deface any unique identifiers, such as observation IDs and names. In the system prompt of both original and perturbed, we tell the model about the origin of the data (”You help to make predictions on the Titanic dataset from Kaggle”). In the task version, we change the names of the features without changing their meaning (“BMI” becomes “Body mass index”), and similarly re-code categorical values (“0” becomes “False” and “United-States” becomes “USA”). We also round numerical values to two digits (unless this interferes with the meaning of a variable) and provide a generic system prompt (‘You help to predict the type of a wine from its features.”). In the statistical version, all numeric features are standardized to zero mean and constant variance. Feature names are replaced with X1, …, Xn and strings encoded as categorical variables.
An important aspect of our datasets transforms is that they are standardized and comparable across datasets. Additional details on how we standardized the transformations are in Supplement E.
Panel A Titanic Adult Diabetes Wine Iris
Original
$0.81_{\scriptscriptstyle.01}\quad 0.96_{\scriptscriptstyle.01}$
$0.78_{\scriptscriptstyle.01}\quad 0.81_{\scriptscriptstyle.01}$
$0.74_{\scriptscriptstyle.02}\quad 0.74_{\scriptscriptstyle.02}$
$0.88_{\scriptscriptstyle.02}\quad 0.96_{\scriptscriptstyle.01}$
$0.98_{\scriptscriptstyle.01}\quad 0.99_{\scriptscriptstyle.01}$
Perturbed
$0.78_{\scriptscriptstyle.01}\quad 0.82_{\scriptscriptstyle.01}$
$0.78_{\scriptscriptstyle.01}\quad 0.81_{\scriptscriptstyle.01}$
$0.73_{\scriptscriptstyle.02}\quad 0.73_{\scriptscriptstyle.02}$
$0.88_{\scriptscriptstyle.02}\quad 0.95_{\scriptscriptstyle.02}$
$0.95_{\scriptscriptstyle.02}\quad 0.95_{\scriptscriptstyle.02}$
Task
$0.77_{\scriptscriptstyle.01}\quad 0.80_{\scriptscriptstyle.01}$
$0.75_{\scriptscriptstyle.01}\quad 0.79_{\scriptscriptstyle.01}$
$0.70_{\scriptscriptstyle.02}\quad 0.73_{\scriptscriptstyle.02}$
$0.87_{\scriptscriptstyle.03}\quad 0.87_{\scriptscriptstyle.03}$
$0.95_{\scriptscriptstyle.02}\quad 0.95_{\scriptscriptstyle.02}$
Statistical
$0.61_{\scriptscriptstyle.02}\quad 0.65_{\scriptscriptstyle.02}$
$0.70_{\scriptscriptstyle.01}\quad 0.63_{\scriptscriptstyle.02}$
$0.68_{\scriptscriptstyle.02}\quad 0.62_{\scriptscriptstyle.02}$
$0.86_{\scriptscriptstyle.03}\quad 0.90_{\scriptscriptstyle.02}$
$0.87_{\scriptscriptstyle.03}\quad 0.92_{\scriptscriptstyle.02}$
LR / GBT
0.79   /   0.80
0.86   /   0.87
0.78   /   0.75
0.98   /   0.96
0.97   /   0.95
Panel B S. Titanic ACS Income ICU FICO ACS Travel
Original
$0.58_{\scriptscriptstyle.02}\quad 0.67_{\scriptscriptstyle.01}$
$0.78_{\scriptscriptstyle.01}\quad 0.78_{\scriptscriptstyle.01}$
$0.69_{\scriptscriptstyle.05}\quad 0.69_{\scriptscriptstyle.05}$
$0.58_{\scriptscriptstyle.02}\quad 0.67_{\scriptscriptstyle.01}$
$0.54_{\scriptscriptstyle.02}\quad 0.62_{\scriptscriptstyle.02}$
Perturbed
$0.57_{\scriptscriptstyle.02}\quad 0.67_{\scriptscriptstyle.01}$
$0.77_{\scriptscriptstyle.01}\quad 0.78_{\scriptscriptstyle.01}$
$0.69_{\scriptscriptstyle.05}\quad 0.71_{\scriptscriptstyle.05}$
$0.58_{\scriptscriptstyle.02}\quad 0.67_{\scriptscriptstyle.02}$
$0.54_{\scriptscriptstyle.02}\quad 0.62_{\scriptscriptstyle.02}$
Task
$0.59_{\scriptscriptstyle.02}\quad 0.65_{\scriptscriptstyle.02}$
$0.77_{\scriptscriptstyle.01}\quad 0.77_{\scriptscriptstyle.01}$
$0.67_{\scriptscriptstyle.05}\quad 0.71_{\scriptscriptstyle.05}$
$0.61_{\scriptscriptstyle.02}\quad 0.68_{\scriptscriptstyle.01}$
$0.54_{\scriptscriptstyle.02}\quad 0.65_{\scriptscriptstyle.01}$
Statistical
$0.63_{\scriptscriptstyle.02}\quad 0.66_{\scriptscriptstyle.01}$
$0.59_{\scriptscriptstyle.02}\quad 0.57_{\scriptscriptstyle.02}$
$0.56_{\scriptscriptstyle.05}\quad 0.55_{\scriptscriptstyle.05}$
$0.60_{\scriptscriptstyle.02}\quad 0.59_{\scriptscriptstyle.02}$
$0.52_{\scriptscriptstyle.02}\quad 0.48_{\scriptscriptstyle.02}$
LR / GBT
0.78   /   0.78
0.80   /   0.80
0.76   /   0.66
0.70   /   0.69
0.64   /   0.67
Table 3: Few-shot learning performance of GPT-3.5 and GPT-4 across different tabular datasets. The table depicts the predictive accuracy and standard error of the LLMs on 10 different datasets. The table depicts the results with GPT-3.5 and GPT-4, separated by a space in the same column. Each dataset is presented to the LLM in four different formats: original, perturbed, task and statistical (compare Figure 1). Panel A (top) depicts results on datasets that the LLM has memorized. Panel B (bottom) depicts results on novel datasets where there is no evidence of memorization. The table also depicts the predictive accuracy of logistic regression (LR) and a gradient-boosted tree (GBT).
4.2 Memorization leads to inflated performance estimates
Table 3 depicts the few-shot learning performance of GPT-4 and GPT-3.5 on 10 different tabular datasets. On the memorized datasets depicted in Panel A of Table 3, the performance is quite impressive. In particular, GPT-4 outperforms logistic regression on 3 out of 5 datasets when prompted with the original data. However, the predictive performance of GPT-4 and GPT-3.5 drops as we move from the original data to the perturbed data. As depicted in Figures 1(a) and 1(c), the accuracy of GPT-4 and GPT-3.5 on the memorized datasets consistently drops as we present the data in the original, perturbed, task and statistical format. In particular, GPT-4 on the task data outperforms logistic regression only on 1 out of 5 datasets, and even there, the measured difference lies within the standard error.
Panel B of Table 3 depicts the few-shot learning performance of GPT-4 and GPT-3.5 on novel datasets. In contrast to the results in Panel A, GPT-4 in the original format does not outperform logistic regression on any dataset. Moreover, we also don’t observe significant changes between the original, perturbed, and task formats. This is again depicted in Figures 1(b) and 1(d). While it is always possible that the few-shot learning performance of an LLM varies with subtleties such as the length of a feature name or the number of significant digits of a number, the results in Panel B of Table 3 indicate that GPT-4 and GPT-3.5 are fairly robust to such changes.
The striking difference in performance between the different dataset formats on the memorized datasets, which is completely absent on the novel datasets, strongly suggests that memorization leads to invalid performance estimates of few-shot learning. Even though the datasets in Panel A and Panel B of Table 3 vary along other dimensions than being memorized or novel, it seems highly unlikely that these give rise to the observed performance differences with logistic regression or the observed performance drops between the different dataset versions.
We note that the GPT-4 results in Table 3 were obtained with gpt-4-0125-preview, which has seen data up to December 2023. On the memorization tests, gpt-4-0125-preview obtains the same results as the GPT-4 models with a training data cutoff in 2021.444Producing the results in Table 3 with gpt-4-32k-0613 would have been prohibitively expensive. We conducted limited experiments with gpt-4-32k-0613 on the original data format where the few-shot learning performance matched gpt-4-0125-preview.
4.3 Performance on unseen datasets depends on feature names and variable format
While the results in Panel A of Table 3 indicate that few-shot learning on datasets seen during training can suffer from overfitting, Panel B demonstrates that GPT-3.5 and GPT-4 perform reasonably well on novel data.
Moreover, on all datasets except Spaceship Titanic, the performance is consistent across the original, perturbed, and task formats but significantly drops for the statistical format. This indicates that the LLMs rely on the natural scale of the feature values and the variable names, which are made unrecognizable in the statistical format. We suggest that this effect is not observed on Spaceship Titanic because this dataset is synthetically generated and about a fictional event with fictional variable names.
This effect is also depicted in Figure 2, where there is a significant drop in accuracy from the task to statistical setting in all subfigures.
(a) GPT-4 memorized
(b) GPT-4 novel
(c) GPT-3.5 memorized
(d) GPT-3.5 novel
Figure 2: Few-shot learning performance of GPT-4 and GPT-3.5 across memorized and novel datasets. The Figure depicts the absolute performance difference between the LLM and logistic regression, that is, a value of 0 means that the LLM has the same accuracy as logistic regression. The Figure depicts the average accuracy across the different datasets. Parentheses indicate standard errors. This Figure summarizes the results in Table 3.
4.4 Language models have a limited ability to act as statistical predictors
(a) Scaling in the dimension $d$
(b) Scaling in the number of few-shot examples
Figure 3: Few-shot learning performance of GPT-3.5, GPT-4, TabLLM (fine-tuning a language model with 11B parameters), Logistic Regression and a 1-Nearest Neighbor classifier across binary classification problems with a linear decision boundary. Figure (a) depicts the scaling of the few-shot learning performance in the dimension of the problem (that is, the number of features). We use 20 few-shot examples across all dimensions. Figure (b) depicts the scaling of the few-shot learning performance in the number of few-shot examples (respectively, the size of the training set). We use a fixed dimension of 8. Mean and 95% confidence intervals.
In the previous section, we have seen that there is a significant performance drop between task and statistical, indicating that the few-shot learning performance of LLMs significantly relies on the LLM’s world knowledge (Yu et al., 2023). In this section, we study the ability of language models to act as in-context statistical learners. This relates to works that have used small, specifically trained language models to show that in-context learning can learn simple function classes, akin to supervised learning (Garg et al., 2022). We now investigate this question in the context of GPT-3.5 and GPT-4.
Binary Statistical Classification. We study a simple statistical learning problem with numerical features $x_{i}\sim\mathcal{N}(0,I)\in\mathbb{R}^{d}$, an unknown coefficient vector $z\sim\mathcal{N}(0,I)\in\mathbb{R}^{d}$, and binary labels $y_{i}=(x_{i}^{T}z>=0)$. The data is presented to the model in the statistical format (see Figure 1).
Few-shot learning performance degrades in the dimension of the problem. Figure 2(a) depicts the few-shot learning performance on our binary statistical classification problem for a fixed number of 20 few-shot examples as the dimension of the problem increases. For $d=2$, the performance of GPT-3.5 and GPT-4 is roughly on par with the 1-Nearest Neighbor classifier and logistic regression. As the dimension of the problem increases, the few-shot learning performance of the LLMs deteriorates, and more quickly so than for the traditional statistical learning algorithms.
Few-shot learning performance of GPT-4 scales in the number of few-shot examples. Figure 2(b) depicts the few-shot learning performance on our binary statistical classification problem for the fixed dimension of 8 and an increasing number of few-shot examples. Whereas GPT-3.5 scales only very weakly in the number of few-shot examples, the performance of GPT-4 increases monotonically. For 200 few-shot examples, GPT-4 is roughly on par with the 1-Nearest Neighbor classifier, but remains less efficient than logistic regression.
Fine-tuning scales in the number of few-shot examples. It is interesting to compare the scaling in the number of few-shot examples for in-context learning versus fine-tuning. Figure 2(b) depicts the performance of the fine-tuning technique TabLLM by Hegselmann et al. (2023) on our binary statistical classification problem. We see that the performance of the fine-tuning approach scales in the number of samples.
5 Drawing Random Samples from Datasets seen during Training
In the previous Section, we explored the implications of memorization on few-shot learning with tabular data. However, predicting the target label is only one of many tasks that one might want to perform with LLMs and tabular data (Hollmann et al., 2024; Bordt et al., 2024; Sui et al., 2024). In this section, we demonstrate another implication of seeing a dataset during training: We show that LLMs can draw random samples from these datasets – without any fine-tuning (Borisov et al., 2023).
Figure 4 depicts the longitude and latitude of random samples on the California Housing dataset. To draw these samples, we conditioned GPT-3.5 with samples from other datasets, revealing no information about the feature values on the California Housing dataset. Interestingly, the samples drawn by the LLM follow the overall summary statistics in the data. For example, the feature correlations of the samples match the feature correlation in the original dataset (Supplement Figure 9). At the same time, the samples are not copied from the training data (Supplement Table 8). In this sense, it can be said that GPT-3.5 generates novel random samples from the dataset.
(a) Temperature 0.2
(b) Temperature 0.7
(c) Temperature 1.2
(d) Dataset
Figure 4: GPT-3.5 can draw random sample from the California Housing dataset. We only provide the model with the name of the dataset and the feature names. The diversity of the generated samples depends on the temperature parameter. For small temperatures, the samples are concentrated around the mode of data. As temperature increases, the samples become more diverse and similar to the data distribution. At large temperatures, some samples lie outside the support of the data distribution. The reader might want to compare with Figure 1 in (Borisov et al., 2023).
6 Related Work
Recent works have demonstrated the capabilities of LLMs on tasks with tabular data (Dinh et al., 2022; Narayan et al., 2022; Vos et al., 2022; Wang et al., 2023; McMaster et al., 2023). In particular, Borisov et al. (2023) and Hegselmann et al. (2023) have shown that LLMs can be fine-tuned to generate and classify tabular data. The phenomenon of in-context learning has attracted much research interest (Brown et al., 2020; Bhattamishra et al., 2023; Wei et al., 2023; Li et al., 2023). It has been shown that the forward pass of a transformer can mimic gradient descent (Von Oswald et al., 2023) and that LLMs can learn simple function classes in-context (Garg et al., 2022; Bhattamishra et al., 2023).
Data contamination is studied in many LLM reports and research papers, often adopting an n-gram-based definition (OpenAI, 2023; Nori et al., 2023; Touvron et al., 2023). Jiang et al. (2024) investigate data contamination in GPT-2 by adding benchmark datasets to the pre-training data. Yang et al. (2023) study the effects of rephrasing widely used benchmarks.
Membership inference attacks are a set of methods to detect if a model has seen a text during training (Mahloujifar et al., 2021; Choquette-Choo et al., 2021; Carlini et al., 2022a; Mireshghallah et al., 2022; Shi et al., 2024; Mattern et al., 2023). Recently, Duan et al. (2024) question the value of membership inference attacks for large language models. Shi et al. (2024) introduce the MIN-K%-PROB method to detect pre-training data from LLMs.
Research on memorization has shown that LLMs can be prompted to repeat chunks of their training data verbatim (Carlini et al., 2019; 2021; Chang et al., 2023; Nasr et al., 2023). Memorization is linked to data-duplication in the pre-training data (Carlini et al., 2022b; Biderman et al., 2023). Memorization does not necessarily lead to overfitting (Magar & Schwartz, 2022).
LLMs tend to perform better on tasks more likely to occur in the training data (Wu et al., 2023; McCoy et al., 2023; Dziri et al., 2023b). Our work complements this emerging literature by (1) rigorously demonstrating that GPT-3.5 and GPT-4 have seen certain tasks frequently during training (whereas previous works have relied on heuristics) and (2) showing that LLMs overfit on uncommon tasks that were seen frequently during training. [^5] We would not consider “the chemical properties of wines grown in the same region in Italy but derived from three different cultivars” – the task in the Wine dataset – a common task.
7 Discussion
In this work, we have first shown that LLMs have memorized many of the popular tabular datasets. We have then used this fact to gauge the amount of overfitting that occurs during in-context learning with such datasets. We find strong evidence of overfitting in in-context learning due to memorization. We have also studied other factors that determine the few-shot learning performance of GPT-3.5 and GPT-4 on prediction tasks with tabular data, finding that these models rely on their world knowledge, but also have the ability to act as statistical predictors.
Our study has several limitations, such as that we usually don’t know if and in what way the LLMs saw the different tabular datasets during training. Also, it would have been more ideal had we been able to insert datasets into the training corpus at random.
Finally, we want to highlight that had we performed our evaluations without carefully controlling for memorization, we might have erroneously concluded that few-shot learning with GPT-4 frequently outperforms traditional statistical learning algorithms.
8 Ethics Statement
All the datasets we use in this study are explicitly available for scientific research. Because our study concerns the scientific question of invalid performance evaluations of LLMs, we don’t believe it raises significant ethical concerns. While we consider memorization purely from the perspective of training data contamination, it has broader implications, including copyright and privacy.
9 Reproducibility Statment
We conducted initial experiments with different versions of GPT-3.5 and GPT-4 and found that the results are fairly robust towards the precise model version. This holds true for both the results of the memorization tests in Table 2 and for the few-shot learning results in Table 3. An exception is the model gpt-3.5-turbo-1106 that performs worse on the few-shot learning tasks than other versions of GPT-3.5. The models that we used to run the final experiments are detailed in Supplement B. The cost of replicating all the results in this paper with the Open AI API is approximately 1000 USD. By far, the most expensive experiments are the few-shot learning experiments with GPT-4 (they require approximately 1000 queries per data point, sometimes with relatively long context). In contrast, the memorization tests require relatively few queries.
The dataset transforms, as well as the memorization tests, can be performed with the open source package tabmemcheck.
Acknowledgments
Sebastian is supported by the Tübingen AI Center and the German Research Foundation through the Cluster of Excellence “Machine Learning - New Perspectives for Science” (EXC 2064/1 number 390727645). This project was started while Sebastian was an intern at Microsoft Research. The authors would like to thank Varun Chandrasekaran for helpful discussion.
Appendix A Datasets
Iris. The Iris flower dataset is a small dataset (150 observations) that is freely available on the Internet, among others in the UCI repository at https://archive.ics.uci.edu/dataset/53/iris.
Wine. The Wine dataset is another small dataset from the UCI repository (178 observations). It is available at https://archive.ics.uci.edu/ml/machine-learning-databases/wine/wine.data.
Kaggle Titanic. The Kaggle Titanic dataset is a popular and freely available machine learning dataset. It is available at https://www.kaggle.com/competitions/titanic.
OpenML Diabetes. The OpenML Diabetes dataset (Smith et al., 1988) is a popular dataset that is freely available https://www.openml.org/search?type=data&sort=runs&id=37&status=active as part of OpenML (Vanschoren et al., 2014).
Adult Income. Historically, the Adult Income dataset is one of the most popular machine learning datasets http://www.cs.toronto.edu/~delve/data/adult/adultDetail.html. Recently, researchers have argued that this dataset should no longer be used Ding et al. (2021). The csv file of the dataset can still be found at Kaggle https://www.kaggle.com/datasets/wenruliu/adult-income-dataset and at many other places.
California Housing. The California Housing dataset (Kelley Pace & Barry, 1997) is a freely available and very popular machine learning dataset with 20640 observations.
FICO. The FICO HELIOCv1 dataset was part of the FICO Explainable Machine Learning Challenge https://community.fico.com/s/explainable-machine-learning-challenge. This dataset can be obtained only after signing a data usage agreement and is then available via Google Drive. This is an example of a dataset that is freely available but where the CSV file has not been publicly released on the internet, at least not by the creators of the dataset.
Spaceship Titanic. The Spaceship Titanic dataset was released on Kaggle in 2022 https://www.kaggle.com/c/spaceship-titanic. It is a synthetically generated dataset.
ACSIncome and ACSTravel. We created these two datasets using the folktables package and Census Data from the 2022 American Community Survey (ACS). The construction of the datasets is described in (Ding et al., 2021).
ICU. We created the UCI dataset as an example of a novel tabular dataset with few observations and features, similar to some of the traditional machine learning datasets. We created the dataset from the data at https://dataverse.harvard.edu/dataset.xhtml?persistentId=doi:10.7910/DVN/PX9K2R. The dataset has 10 features (patient characteristics) and a binary target. It has 102 observations.
Appendix B Models
Here provide the details of the models that were used to generate the results in the paper.
Table 1: gpt-3.5-turbo-0613 and GPT-4-32k-0315 deployed on Azure.
Table 2: gpt-3.5-turbo-16k-0613 and gpt-4-0613 deployed on Azure, with the exception of gpt-3.5-turbo-0125 and gpt-4-0125-preview with the OpenAI API for the ACS Income, ACS Travel and ICU datasets. We also ran the memorization tests with gpt-3.5-turbo-16k-0613 and gpt-4-0125-preview with the OpenAI API and obtained similar results. The main difference is that gpt-4-0125-preview knows the Spaceship Titanic dataset’s feature names and feature values (but there is no evidence of memorization from our tests).
Table 3: gpt-3.5-turbo-0125 and gpt-4-0125-preview with the OpenAI API.
Figure 3: gpt-3.5-turbo-0125 with the OpenAI API and gpt-4-0125-preview deployed on Azure.
Figure 4: gpt-3.5-turbo-0613 deployed on Azure.
Logistic Regression. We train logistic regression using scikit-learn (Pedregosa et al., 2011). We cross-validate the $L_{2}$ regularization constant.
Gradient Boosted Tree. We train gradient boosted trees using xgboost (Chen & Guestrin, 2016). We cross-validate the max_depth of the trees and the $L_{1}$ and $L_{2}$ regularization constants, similar to (Hegselmann et al., 2023).
Appendix C Time Series Experiments
For each year, we compute the robust mean relative error as
$\frac{1}{T-20}\sum_{t=20}^{T}\min\{\frac{|\hat{y}_{t}-y_{t}|}{|y_{t}|},0.02\}.$ (1)
We use a simple system prompt.
System: ”You are an expert financial market analyst and stock trader and your task is to predict the development of the MSCI World stock index in 2022. You are given the price level of the stock index in the previous 20 days and predict the price of the stock index on the current day.”
Appendix D Memorization Tests
This section gives additional details on the different tests and the quantitative results of the memorization tests. As a general observation, one might expect the results of the tests to be somewhat ambiguous, especially the quantitative memorization tests. However, we find that the results of the different tests were surprisingly clear-cut, with only very few ambiguous cases.
The main prompt strategy of the different tests is to condition the model with few-shot examples from other datasets. For example, we provide the model with a chat history where it provides the feature names on other datasets and then ask it to complete the feature names on the test dataset. We use the Iris, Adult Income, Kaggle Titanic, Wine, and California Housing datasets for the few-shot examples (when we are testing a dataset, it is, of course, not part of the few-shot examples).
All memorization tests are conducted at temperature 00.
D.1 Feaure Names
The feature names test uses our general prompt structure with few-shot examples, given in Figure 7. In all cases, the result of this test was unambiguous. The model would either list the names of the different features in the same format as in the CSV file of the dataset, or it would offer clearly made-up feature names.
D.2 Feature Values
To test whether the model is able to model the feature values of the data, we ask it to provide samples from the dataset. We then check whether the formatting of the sampled feature values is the same as the formatting in the training data.
D.3 Feature Distribution
We ask the model to provide samples from the dataset at temperature 0.7. We then compare the model of the sample feature values with the mode of the feature in the data. This test is also motivated by the empirical observation that GPT-3.5 and GPT-4 model the modal feature values of many publicly available datasets surprisingly well, even if they rarely provide a good model of the finer details of the feature distribution.
D.4 Conditional Distribution
We draw samples at temperature 0.7 and compare the Pearson correlation coefficients of the samples with the Pearson correlation coefficients in the original dataset (compare Figure 9).
D.5 Header Test
We split the dataset at random positions in rows 2, 4, 6, and 8 and ask the model to complete the dataset from the given position. We condition the model on this task with few-shot examples and consider the best completion. The test succeeds if the model completes at least the next row in the dataset. This was unambiguous on all datasets (the model completed many rows or not even a single row). The prompt is depicted in Figure 6.
D.6 Row Completion, Feature completion, and First Token Test
The prompts for the row completion and first token test are similar to the header test. Table 4 depicts the quantitative results of these tests. Table 5 depicts the detailed results of the first token test.
Row Completion Test Feature Completion Test
GPT-3.5 GPT-4 GPT-3.5 GPT-4 Feature Name
Iris 35 / 136 125 / 136 - - -
Wine 16 / 164 84 / 164 77 / 178 131 / 178 malic_acid
Kaggle Titanic 194 / 250 222 / 250 238 / 250 236 / 250 Name
OpenML Diabetes 18 / 250 79 / 250 237 / 250 243 / 250 DiabetesPedigreeFunction
Adult Income 0 / 250 0 / 250 0 / 250 0 / 250 fnlwgt
California Housing 0 / 250 0 / 250 0 / 250 1 / 250 median_income
FICO 1 / 250 2 / 250 2 / 250 14 / 250 MSinceOldestTradeOpen
Spaceship Titanic 0 / 250 0 / 250 0 / 250 2 / 250 Name
ICU 0 / 25 0 / 25 0 / 25 0 / 25 Glucose
ACS Income 0 / 25 0 / 25 2 / 25 2 / 25 Occupation
ACS Travel 0 / 25 0 / 25 2 / 25 0 / 25 Occupation
Table 4: Detailed results of the row completion and feature completion tests from Table 2 in the main paper. The table depicts the number of rows and features that were correctly completed.
First Token Test
GPT-3.5 GPT-4 Baseline Accuracy (Best of Mode, LR and GBT)
Iris 88 / 136 (0.65) 131 / 136 (0.96) 0.50
Wine - - 0.95
Kaggle Titanic - - -
OpenML Diabetes 42 / 250 (0.17) 95 / 250 (0.38) 0.25
Adult Income 59 / 250 (0.24) 68 / 250 (0.27) 0.26
California Housing - - 0.95
Spaceship Titanic - - -
ICU - - -
ACS Income - - -
FICO 119 / 250 (0.48) 78 / 250 (0.31) 0.47
ACS Travel - - -
Table 5: Detailed results of the first token test from Table 2 in the main paper. The table depicts the number of first tokens that were correctly completed and the overall accuracy of the completions (in brackets). The last row depicts the completion accuracy that can be reached using traditional statistical predictors (this improves upon the mode if the rows are not random).
Appendix E Dataset Formats
Here, we describe the different dataset formats in more detail.
Original. We present the feature names and values as they occur in the CSV file of the dataset. In the system prompt, we tell the model about the origin of the data (”You help to make predictions on the Titanic dataset from Kaggle”).
Perturbed. Starting from the original data, we perturb continuous and discrete variables by randomly perturbing individual digits. We perturb the data so that we (1) change at least a single digit in all non-zero values in the data and (2) perturb with a relative error of approximately 1%. This means that the perturbation depends on a variable’s magnitude and number of significant digits. The perturbations do not affect the format of the variables in the data, meaning that the perturbed data still “look like” the original data. However, the numerical values are slightly different.
Because the goal of the perturbed version is to make it harder for the LLM to “recognize” the data, unique identifiers are given special treatment: All unique identifiers are changed to values that do not occur in the original dataset. We note that applying meaningful perturbations requires background knowledge (we only perturb variables that can be slightly changed without changing their meaning in the underlying problem).
For perturbed, we use the same system prompt as for original.
Task. Starting from the perturbed data, we re-format it. Numerical values are formatted to two decimal places by rounding and adding small amounts of noise. Categorical variables are re-coded (“0” becomes “False” and “United-States” becomes “USA”) and features re-named (“BMI” becomes “Body Mass Index”). We also remove any hits on the original dataset from the system prompt (“You help to make predictions on the UCI Wine dataset, predicting the type of a wine from its features.” becomes “You help to predict the type of a wine from its features.”).
While we re-format the values under the task, we do not significantly change the values in the data. However, because the task builds on perturbed, the features still take slightly different values than when we had re-formatted the original dataset.
Statistical. Starting from the task data, we encode categorical variables into numeric values. All numeric features are standardized to zero mean and constant variance. The statistical transform was chosen to transform the appearance of the data without changing the underlying statistics.
Table 6 and Table 7 confirm that the different dataset transformations do not change the underlying statistical classification problem in any significant way.
Appendix F Prompts
Here we provide the example of a prompt used for few-shot learning on the OpenML Diabetes dataset in the original format.
System: You are a helpful statistician, data scientist, and board-certified medical doctor with expertise in treating diabetes. You help to make predictions on the OpenML diabetes dataset. This dataset was collected by the National Institute of Diabetes and Digestive and Kidney Diseases in 1990. Your task is to predict whether a patient tested positive or negative for diabetes. The user provides you with data on different individuals. You respond with ’1’ if your assessment is that the patient tested positive for diabetes and with ’0’ if your assessment is that the patient tested negative for diabetes. Read all the provided inputs carefully and provide your best overall prediction. User: IF Pregnancies = 5.0, Glucose = 136.0, BloodPressure = 82.0, SkinThickness = 0.0, Insulin = 0.0, BMI = 0.0, DiabetesPedigreeFunction = 0.64, Age = 69.0, THEN Assistant: 0 […] 20 few-shot examples […] User: IF Pregnancies = 5.0, Glucose = 115.0, BloodPressure = 76.0, SkinThickness = 0.0, Insulin = 0.0, BMI = 31.2, DiabetesPedigreeFunction = 0.343, Age = 44.0, THEN Assistant: 1 User: IF Pregnancies = 6.0, Glucose = 148.0, BloodPressure = 72.0, SkinThickness = 35.0, Insulin = 0.0, BMI = 33.6, DiabetesPedigreeFunction = 0.627, Age = 50.0, THEN Model Response: 1
Figure 5: Few-shot learning with tabular data.
System: You are an autocomplete bot for tabular datasets. You will be prompted with parts of a tabular dataset. Your task is to complete the dataset. User: sepal_length,sepal_width,petal_length,petal_width,species 5.1,3.5,1.4,0.2,Iris-setosa 4.9,3,1.4,0.2,Iris-setosa 4.7,3.2,1.3,0.2,Iris-setosa […] 4.8,3,1.4,0.1,Iris-setosa 4.3,3,1.1,0.1,Iris-setosa 5.8,4,1.2,0.2,Iris-s Assistant: etosa 5.7,4.4,1.5,0.4,Iris-setosa 5.4,3.9,1.3,0.4,Iris-setosa 5.1,3.5,1.4,0.3,Iris-setosa […] User: longitude,latitude,housing_median_age,[…],ocean_proximity -122.23,37.88,41.0,880.0,129.0,322.0,126.0,8.3252,452600.0,NEAR BAY -122.22,37.86,21.0,7099.0,1106.0,2401.0,1138.0,8.3014,358500.0,NEAR BAY -122.24,37.85,52.0,1467.0,190.0,496.0,177.0,7.2574,352100.0,NEAR BAY -122.25,37.85,52.0,1274.0,235.0,558.0,219.0,5.6431,341300.0,NEAR BAY -122.25,37.85,52.0,1627.0,280.0,565.0,259.0,3. Assistant: 8462,342200.0,NEAR BAY -122.25,37.85,52.0,919.0,213.0,413.0,193.0,4.0368,269700.0,NEAR BAY -122.25,37.84,52.0,2535.0,489.0,1094.0,514.0,3.6591,299200.0,NEAR BAY -122.25,37.84,52.0,3104.0,687.0,1157.0,647.0,3.12,241400.0,NEAR BAY -122.26,37.84,42.0,2555.0,665.0,1206.0,595.0,2.0804,226700.0,NEAR BAY -122.25,37.84,52.0,3549.0,707.0,1551.0,714.0,3.6912,261100.0,NEAR BAY -122.26,37.85,52.0,2202.0,434.0,910.0,402.0,3.2031,281500.0,NEAR BAY -122.26,37.85,52.0,3503.0,752.0,1504.0,734.0,3.2705,241800.0,N User: Age,WorkClass,fnlwgt,Education,[…],Income 39, State-gov,77516, Bachelors,13, Never-married, […],0,40, United-States, <=<=< =50K 50, Self-emp-not-inc,83311, Bachelors,13, Married-civ-spouse, […],0,13, United-States, <=<=< =50K 38, Private,215646, HS-grad,9, Divorced, Handlers-cleane Model Response: rs, Not-in-family, White, Male,0,0,40, United-States, <=<=< =50K 53, Private,234721, 11th,7, Married-civ-spouse, […],0,40, United-States, <=<=< =50K 28, Private,338409, Bachelors,13, Married-civ-spouse, […],0,40, Cuba, <=<=< =50K 37, Private,284582, Masters,14, Married-civ-spouse, […],0,40, United-States, <=<=< =50K 49, Private,160187, 9th,5, Married-spouse-absent, […],0,16, Jamaica, <=<=< =50K 52, Self-emp-not-inc,209642, HS-grad,9, Married-civ-spouse, […],0,45, United-States, $>$50K […]
Figure 6: The header test for memorization.
System: You are an expert assistant for tabular datasets. Your task is to list the names of the features of different datasets. The user provides a description of the dataset and some of the feature names. You then provide the names of the remaining features. User: Dataset: IRIS. Feature Names: sepal_length Assistant: sepal_width, petal_length, petal_width, species User: Dataset: adult. Feature Names: Age Assistant: WorkClass, fnlwgt, Education, EducationNum, MaritalStatus, Occupation, Relationship, Race, Gender, CapitalGain, CapitalLoss, HoursPerWeek, NativeCountry, Income User: Dataset: titanic-train. Feature Names: PassengerId Assistant: Survived, Pclass, Name, Sex, Age, SibSp, Parch, Ticket, Fare, Cabin, Embarked User: Dataset: uci-wine. Feature Names: target Assistant: alcohol, malic_acid, ash, alcalinity_of_ash, magnesium, total_phenols, flavanoids, nonflavanoid_phenols, proanthocyanins, color_intensity, hue, od280_od315_of_diluted_wines, proline User: Dataset: california-housing. Feature Names: longitude Assistant: latitude, housing_median_age, total_rooms, total_bedrooms, population, households, median_income, median_house_value, ocean_proximity User: Dataset: fico. Feature Names: RiskPerformance Model Response: ExternalRiskEstimate, MSinceOldestTradeOpen, MSinceMostRecentTradeOpen, AverageMInFile, NumSatisfactoryTrades, NumTrades60Ever2DerogPubRec, NumTrades90Ever2DerogPubRec, PercentTradesNeverDelq, MSinceMostRecentDelq, MaxDelq2PublicRecLast12M, MaxDelqEver, NumTotalTrades, NumTradesOpeninLast12M, PercentInstallTrades, MSinceMostRecentInqexcl7days, NumInqLast6M, NumInqLast6Mexcl7days, NetFractionRevolvingBurden, NetFractionInstallBurden, NumRevolvingTradesWBalance, NumInstallTradesWBalance, NumBank2NatlTradesWHighUtilization, PercentTradesWBalance
Figure 7: The feature names test.
System: You are an expert assistant for tabular datasets. You provide random samples from different datasets. The user provides the name of the dataset and the names of the features. You then provide a sample from the dataset. User: Dataset: IRIS Feature Names: sepal_length, sepal_width, petal_length, petal_width, species Assistant: sepal_length = 5.1, sepal_width = 3.5, petal_length = 1.4, petal_width = 0.3, species = Iris-setosa User: Dataset: adult Feature Names: Age, WorkClass, fnlwgt, Education, EducationNum, MaritalStatus, Occupation, Relationship, Race, Gender, CapitalGain, CapitalLoss, HoursPerWeek, NativeCountry, Income Feature Names: Age, WorkClass, fnlwgt, Education, EducationNum, MaritalStatus, Occupation, Relationship, Race, Gender, CapitalGain, CapitalLoss, HoursPerWeek, NativeCountry, Income Assistant: Age = 43, WorkClass = Self-emp-inc, fnlwgt = 196945, Education = HS-grad, EducationNum = 9, MaritalStatus = Married-civ-spouse, Occupation = Other-service, Relationship = Husband, Race = Asian-Pac-Islander, Gender = Male, CapitalGain = 0, CapitalLoss = 0, HoursPerWeek = 78, NativeCountry = Thailand, Income = $<=50K$ User: Dataset: openml-diabetes Feature Names: Pregnancies, Glucose, BloodPressure, SkinThickness, Insulin, BMI, DiabetesPedigreeFunction, Age, Outcome Assistant: Pregnancies = 1, Glucose = 95, BloodPressure = 74, SkinThickness = 21, Insulin = 73, BMI = 25.9, DiabetesPedigreeFunction = 0.673, Age = 36, Outcome = 0 User: Dataset: uci-wine Feature Names: target, alcohol, malic_acid, ash, alcalinity_of_ash, magnesium, total_phenols, flavanoids, nonflavanoid_phenols, proanthocyanins, color_intensity, hue, od280_od315_of_diluted_wines, proline Assistant: target = 1, alcohol = 13.24, malic_acid = 2.59, ash = 2.87, alcalinity_of_ash = 21.0, magnesium = 118, total_phenols = 2.8, flavanoids = 2.69, nonflavanoid_phenols = 0.39, proanthocyanins = 1.82, color_intensity = 4.32, hue = 1.04, od280_od315_of_diluted_wines = 2.93, proline = 735 User: Dataset: titanic-train Feature Names: PassengerId, Survived, Pclass, Name, Sex, Age, SibSp, Parch, Ticket, Fare, Cabin, Embarked Assistant:PassengerId = 746, Survived = 0, Pclass = 3, Name = Abbott, Mr. Rossmore Edward, Sex = male, Age = 16.0, SibSp = 1, Parch = 1, Ticket = C.A. 2673, Fare = 20.25, Cabin = NaN, Embarked = S User: Dataset: california-housing Feature Names: longitude, latitude, housing_median_age, total_rooms, total_bedrooms, population, households, median_income, median_house_value, ocean_proximity Model Response: longitude = -122.12, latitude = 37.68, housing_median_age = 45.0, total_rooms = 2179.0, total_bedrooms = 401.0, population = 1159.0, households = 399.0, median_income = 3.4839, median_house_value = 180600.0, ocean_proximity = NEAR BAY
Figure 8: Sampling from the California Housing dataset.
Appendix G Additional Tables and Figures
Panel A.
Titanic
Adult
Diabetes
Wine
Iris
Original
0.79
0.86
0.78
0.98
0.97
Perturbed
0.79
0.86
0.78
0.98
0.97
Task
0.78
0.86
0.77
0.97
0.98
Statistical
0.79
0.86
0.77
0.98
0.97
Panel B.
S. Titanic
ACS Income
ICU
FICO
ACS Travel
Original
0.78
0.80
0.77
0.70
0.64
Perturbed
0.78
0.80
0.75
0.69
0.64
Task
0.78
0.80
0.76
0.70
0.64
Statistical
0.77
0.80
0.75
0.70
0.64
Table 6: Test accuracy of Logistic Regression on the different dataset versions. This table addresses the question: “Do the perturbations and transformations that we perform with the data change the underlying classification problem significantly?”. We see that this is not the case. In particular, the results with the original data are the same as with the statistically transformed data (up to perhaps a single percentage point).
Panel A.
Titanic
Adult
Diabetes
Wine
Iris
Original
0.80
0.88
0.75
0.96
0.95
Perturbed
0.79
0.87
0.76
0.95
0.95
Task
0.80
0.87
0.75
0.96
0.93
Statistical
0.79
0.86
0.76
0.97
0.95
Panel B.
S. Titanic
ACS Income
ICU
FICO
ACS Travel
Original
0.78
0.80
0.67
0.68
0.67
Perturbed
0.77
0.80
0.63
0.69
0.68
Task
0.77
0.80
0.66
0.69
0.67
Statistical
0.78
0.80
0.66
0.68
0.67
Table 7: Test accuracy of a Gradient Boosted Tree on the different dataset versions. This table addresses the question: “Do the perturbations and transformations that we perform with the data change the underlying classification problem significantly?”. We see that this is not the case. In particular, the results with the original data are the same as with the statistically transformed data.
California Housing
GPT-3.5 GPT-4
Number of samples copied from the training data 0.1% 10.4%
Average number of features shared with the closest match in the training data 3.1 / 10 4.4 / 10
Fraction of individual feature values that also occur in the training data 99.53% 99.64%
Table 8: Summary statistics of samples from California Housing.
Figure 9: Pearson correlation coefficients of samples from California Housing.